# Finding Kissing Numbers with Game-theoretic Reinforcement Learning

### v2

@@ Abstract
Since Isaac Newton first studied the Kissing Number Problem in 1694, determining the maximal number of non-overlapping spheres around a central sphere has remained a fundamental challenge. This problem represents the local analogue of Hilbert’s 18th problem on sphere packing, bridging geometry, number theory, and information theory. Although significant progress has been made through lattices and codes, the irregularities of high-dimensional geometry and exponentially growing combinatorial complexity beyond 8 dimensions, which exceeds the complexity of Go game, limit the scalability of existing methods. Here we model this problem as a two-player matrix completion game that can be fully parallelized at large scale, and train the game-theoretic reinforcement learning system, PackingStar, to efficiently explore high-dimensional spaces. The matrix entries represent pairwise cosines of sphere center vectors; one player fills entries while another corrects suboptimal ones, jointly maximizing the matrix size, corresponding to the kissing number. This cooperative dynamics substantially improves sample quality, making the extremely large spaces tractable. PackingStar reproduces previous configurations and surpasses all human-known records from dimensions 25 to 31, with the configuration in 25 dimensions geometrically corresponding to the Leech lattice and suggesting possible optimality. It achieves the first breakthrough beyond rational structures from 1971 in 13 dimensions, discovers over 6000 new structures in 14 and other dimensions, and sets new records for generalized kissing configurations under various angular constraints. These results demonstrate AI’s power to explore high-dimensional spaces beyond human intuition and open new pathways for the Kissing Number Problem and broader geometry problems.
@@ Introduction
The Kissing Number Problem was first posed in 1694 by Isaac Newton and David Gregory, asking for the maximal number of non-overlapping spheres that can touch a central sphere in three-dimensional space. It was later extended to $n$-dimensional Euclidean space, where $K(n)$ denotes the kissing number in $n$ dimensions, sparking centuries of research in high-dimensional geometry. In 1900, David Hilbert included the sphere packing problem in $n$-dimensional space as the 18th problem of his famous 23 problems [?]. The Kissing Number Problem can be viewed as a local analogue of Hilbert’s 18th problem [?]. Beyond geometry, it connects to number theory, group theory, and information theory [?,?,?]. Therefore, advancing this problem could foster progress across multiple fields.
@@
Unfortunately, determining kissing numbers in higher dimensions remains exceptionally difficult. Table 1 lists the kissing numbers in different dimensions. After centuries of effort, the problem has been solved exactly only in a few dimensions 1, 2, 3, 4, 8, and 24 [?,?,?,?,?,?,?]. As the dimension increases, human understanding becomes more limited due to the growing geometric complexity of high-dimensional spaces and combinatorial explosion of configurations. Over the past half-century, mathematicians have continued to push the boundaries using lattice constructions [?], code techniques [?], group representation theory [?], and carefully designed substructures derived from known configurations [?]. In dimensions 10, 11, 13, 14, and dimensions from 17 to 21, such approaches established new lower bounds [?,?]. In dimensions 25–31, studies conducted in 2011 and 2016 further improved the bounds by exploiting subsets of Leech lattice vectors combined with heuristic optimization [?,?]. While these results mark significant milestones, they are limited by reliance on prior knowledge, and painstaking human design. As dimensionality increases, the exponentially growing configuration space and the breakdown of full symmetry assumptions make existing approaches insufficient. These challenges highlight the need for novel paradigms to explore complex structures beyond human intuition.
@@
In this paper, we introduce PackingStar, a multi-agent reinforcement learning system that reformulates the Kissing Number Problem as a two-player matrix completion game. This approach scales efficiently across dimensions, mitigates the combinatorial explosion of high-dimensional search spaces, and overcomes the challenge of extreme data scarcity due to the limited human knowledge of high-dimensional constructions. In PackingStar, two cooperative agents learn to play the matrix completion game to efficiently explore irregular structures in high-dimensional spaces without human prior knowledge. The first agent focuses on incrementally filling matrix entries to construct larger and denser configurations. However, in the early stages of exploration, the candidate space for potential entries is exponentially large, and the agent only observes local matrix information, making erroneous placements inevitable. To address this problem, the second agent is introduced to learn how to identify and remove invalid or inconsistent entries. By leveraging a more global understanding of the matrix geometry and high-dimensional structure, this agent dynamically corrects the mistakes made by the first agent, enabling the joint system to converge toward larger configurations—corresponding to higher kissing numbers. The key insight to PackingStar’s success lies in two complementary design principles. First, the framework reformulates the kissing number construction as an equivalent matrix completion problem in the Gram matrix domain. This reformulation eliminates numerical instability and inefficiency inherent in high-dimensional coordinate representations, enabling all computations to operate directly on the Gram matrix in a fully parallelizable manner. The formulation seamlessly couples problem modeling with its learning methods, allowing efficient GPU acceleration and large-scale sampling for AI-driven discovery of kissing numbers. Second, PackingStar features the finely designed game-theoretic interaction between the two agents. By strategically performing delayed and selective corrections, the second agent imposes geometric constraints on the first agent’s expansive search, effectively regularizing exploration in an otherwise intractable space. This game dynamics not only improves the quality of sampled configurations and the efficiency of value estimation but also reduces the search complexity in exponentially large spaces, enabling scalable discovery of complex high-dimensional configurations that were previously inaccessible to existing methods.
@@
Figure 1: AI breakthroughs in the Kissing Number Problem. Grey markers show the best previously known lower bounds. The blue triangle marks the 11D improvement achieved by AlphaEvolve (2025) [?]. In 13D, PackingStar (red star) uncovers a new rational configuration with kissing number $K_{r}(13)=1146$, distinct from the current best non-rational lower bound $K(13)=1154$. For $25\leq n\leq 31$, the red stars indicate new lower bounds $K_{\mathrm{new}}(n)$ discovered by PackingStar, with the inset reporting the corresponding increments $\Delta K(n)=K_{\mathrm{new}}(n)-K_{\mathrm{prev}}(n)$ compared to the previous best $K_{\mathrm{prev}}(n)$.
@@
PackingStar demonstrates remarkable scalability, successfully reproducing all known kissing number configurations in dimensions up to 31. Figure 1 illustrates some of PackingStar’s breakthrough discoveries on this problem. From dimensions 25 to 31, PackingStar surpasses the lower bounds reported in 2016 and also exceeds the meta-construction form introduced in 2011 [?,?], achieving what appears to be the optimal construction form along this path; the resulting construction form are summarized in Table 2. Notably, the breakthrough configuration discovered in 25 dimensions exhibits a clear geometric pattern that corresponds closely to the substructure of Leech lattice. Across multiple metrics, this configuration shows strong convergence, suggesting that it may represent the optimal 25-dimensional arrangement, although a rigorous mathematical proof is still lacking. In 13 dimensions, PackingStar surpasses the previously known rational structure with 1130 spheres from 1971 [?], reaching a kissing number of 1146. Rational structures are those in which all pairwise cosines between sphere centers are rational. The non-rational structure with 1154 spheres is obtained by slightly adjusting the coordinates of some spheres in the 1130-sphere configuration, later reached 1154 in 1999 [?]. Although the new 13-dimensional rational configuration does not exceed the non-rational record, it is of considerable mathematical value. Its fully rational inter-sphere cosines allow for exact analytic study without relying on numerical approximations. This offers insights into the effects of different constraints on kissing numbers and may inspire higher-density packings. Rational structures also facilitate more efficient encoding and decoding in information and communication applications, highlighting their practical significance alongside their theoretical and geometric value. Beyond these specific dimensions, PackingStar uncovers over 6000 new configurations in 14 dimensions and identifies additional configurations corresponding to current lower bounds in dimensions 12, 13, 15, and others. The sheer volume of these newly discovered arrangements provides rich opportunities for further investigation and may ultimately lead to configurations with even higher kissing numbers. In addition, PackingStar achieves significant new records for generalized kissing configurations under tighter angular constraints, including a 12-dimensional configuration with 81 spheres under the $1/4$ cosine constraint, as well as 14-dimensional and 17-dimensional configurations with 252 and 578 spheres under the $1/3$ constraint, respectively. All of these results have been independently validated and are now included in the standard databases for kissing numbers and spherical codes. In summary, these results highlight AI’s ability to understand and reason about high-dimensional geometric structures, opening a new paradigm for studying the Kissing Number Problem and demonstrating the potential of AI to explore a broader class of challenging mathematical problems.
Key difference to AlphaEvolve
The key difference between PackingStar and AlphaEvolve lies in the search space. AlphaEvolve optimizes explicit coordinates in $\mathbb{R}^{n}$, which is numerically unstable and scales exponentially with dimension; increasing computational resources cannot overcome this fundamental limitation. In contrast, PackingStar operates in the Gram-matrix domain with strict geometric constraints, enabling efficient large-scale sampling. This fundamental distinction explains why PackingStar scales to higher dimensions while AlphaEvolve cannot [?].
Although AlphaEvolve improved the kissing number in 11 dimensions from 592 to 593, likely by adjusting the coordinates of the known 592-sphere configuration, the result lacks clear geometric structure and does not generalize to other dimensions. In contrast, PackingStar not only sets new bounds across higher dimensions but also discovers geometrically meaningful, analyzable configurations, providing concrete objects for further mathematical study [?].
@@ Results
The second objective is to identify improved meta construction forms that assemble these subsets into full configurations. Both objectives are jointly addressed within the unified PackingStar framework, which simultaneously increases the achievable size of $S_{i}$ and uncovers superior meta construction forms. Our findings show that PackingStar not only discovers larger subsets but also identifies a strictly better construction template than the meta-construction introduced in 2011; moreover, evidence from convergence and geometric consistency strongly suggests that the new template is optimal within this construction pathway. These advances yield especially pronounced improvements in dimensions 29 and 31. Table 2 reports additional new construction forms that deviate from the classical form [?,?]. In dimensions 26263030, the system consistently augments the Leech-lattice component by adding lower-dimensional parts $K(k)$ where k=n24k=n-24 (e.g., adding $K(2)$ in dimension $26$). In 31 dimensions, PackingStar uncovers a novel assembly pattern by partitioning a 7-dimensional kissing configuration into 42 (optimal) disjoint unit-radius equilateral triangles, resulting in an $84$-fold weighted $S_{i}$, surpassing the previous $75$-fold weighted $S_{i}$. Similarly, in 29 dimensions, PackingStar embeds 12 (optimal) disjoint unit-radius equilateral triangles into the 5-dimensional kissing configuration to generate a $26$-fold weighted $S_{i}$, exceeding the prior $24$-fold weighted $S_{i}$. These advances yield strictly larger kissing configurations in all dimensions from $25$ to $31$. The resulting new lower bounds $K_{\mathrm{new}}(n)$ are summarized in Table 2. Compared with the previous best values $K_{\mathrm{prev}}(n)$, our constructions increase the kissing number by $8,38,68,152,1224,456,$ and $5476$ spheres in dimensions $25$ through $31$, respectively. These improvements are also highlighted by the red markers in Fig. 1. Table 3 summarizes the cosine value among sphere pairs in the new configurations.
@@
Diverse constructions in different dimensions
New records for generalized kissing configurations
Table 4: New maximal values $K_{\mathrm{new}}(n,\alpha)$ for generalized kissing configurations with maximal inner product at most $\alpha$, obtained by PackingStar, compared with the previous best known values $K_{\mathrm{prev}}(n,\alpha)$. Bold entries indicate improved configurations discovered by PackingStar.
12 $1/4$ $\{-\tfrac{1}{2},\ -\tfrac{1}{8},\ \tfrac{1}{4}\}$ 81 79
As shown in Table 4, the new constructions from PackingStar yield new record for $K(n,\alpha)$ in several dimensions. In 12 dimensions under the $1/4$ cosine constraint, we achieve $K_{\mathrm{new}}(12,1/4)=81$, improving over the previous best of 79. Similarly, in 14 and 17 dimensions under the $1/3$ constraint, we obtain $K_{\mathrm{new}}(14,1/3)=252$ and $K_{\mathrm{new}}(17,1/3)=578$, surpassing the previously reported values of 240 and 546, respectively.
Notably, the $K_{\mathrm{new}}(12,1/4)$ configuration contains no antipodal pairs; however, each of the 81 spheres has the same number of neighbors at each distance, forming a highly regular local structure. Such a combination of high local regularity without any antipodal pair appears to be rare in spherical code. We find several distinct variants of $K_{\mathrm{new}}(12,1/4)$ configuration, among which the most symmetric realization turns out to be the Kronecker product of the Schläfli spherical code with the unit equilateral triangle. This arrangement possesses a transitive symmetry group of order 311040.
In dimension 17, PackingStar found a significant improvement: a configuration with 578 spheres, exceeding the previously known 546. All three geometrically non-isomorphic variants discovered share the same cosine set. Analysis of these variants reveals a remarkable structural property: the $K_{\mathrm{new}}(17,1/3)$ configuration can be interpreted as essentially two copies of the 16-dimensional 288-sphere configuration, which is known to be optimal under the cosine $1/4$ constraint, together with two additional points, i.e., $578=288\times 2+2$. This decomposition provides strong evidence that 578 is very likely optimal for the cosine $1/3$ constraint in dimension 17.

### v3

@@ Abstract
Since Isaac Newton first studied the Kissing Number Problem in 1694, determining the maximal number of non-overlapping spheres around a central sphere has remained a fundamental challenge. This problem is the local analogue of Hilbert’s 18th problem, bridging geometry, number theory, and information theory. Although significant progress has been made through lattices and codes, the irregularities of high-dimensional geometry, dimensional structure variability, and combinatorial explosion beyond Go limit the scalability and generality of existing methods. Here we model the problem as a two-player matrix completion game and train the reinforcement learning system, PackingStar, to play the games. The matrix entries represent pairwise cosines of sphere center vectors. One player fills entries while another corrects suboptimal ones to improve exploration quality, cooperatively maximizing the matrix size, corresponding to the kissing number. These matrices are decomposed into representative substructures, providing diverse bases and structural constraints that steer subsequent games and make extremely large spaces tractable. PackingStar surpasses records from dimensions 25 to 31 and sets new lower bounds for generalized kissing numbers under various angular constraints. It achieves the first breakthrough beyond rational structures from 1971 in 13 dimensions and discovers over 6000 new structures in other dimensions. Notably, some configurations challenge long-held antipodal paradigms, revealing algebraic correspondences with finite simple groups as well as geometric relationships across dimensions. Inspired by these patterns, humans devised further improved constructions. These results demonstrate AI’s power to explore high-dimensional spaces beyond human intuition via extreme-scale reinforcement learning and open new pathways for the Kissing Number Problem and broader geometry research.
@@ Introduction
The Kissing Number Problem was first posed in 1694 by Isaac Newton and David Gregory, asking for the maximal number of non-overlapping spheres that can touch a central sphere in three-dimensional space. It was later extended to $n$-dimensional Euclidean space, where $K(n)$ denotes the kissing number in $n$ dimensions, sparking centuries of research in high-dimensional geometry. In 1900, David Hilbert included the sphere packing problem in $n$-dimensional space as the 18th problem of his famous 23 problems [?]. The Kissing Number Problem is the local analogue of Hilbert’s 18th problem [?]. It connects to geometry, number theory, group theory, and information theory [?,?,?]. Therefore, advancing this problem could foster progress across multiple fields.
@@
Unfortunately, determining kissing numbers in higher dimensions remains exceptionally difficult. Table 1 lists the kissing numbers in different dimensions. After centuries of effort, the problem has been solved exactly only in a few dimensions 1, 2, 3, 4, 8, and 24 [?,?,?,?,?,?,?]. As the dimension increases, human understanding becomes more limited due to the growing geometric complexity of high-dimensional spaces and combinatorial explosion of configurations. Over the past half-century, mathematicians have continued to push the boundaries using lattice constructions [?], code techniques [?], group representation theory [?], and carefully designed substructures derived from known configurations [?]. In dimensions 10, 11, 13, 14, and dimensions from 17 to 21, such approaches established new lower bounds [?,?]. In dimensions 25–31, studies conducted in 2011 and 2016 further improved the bounds by exploiting subsets of Leech lattice vectors combined with heuristic optimization [?,?]. While these results mark significant milestones, they are limited by reliance on prior knowledge, and specific human design. As dimensionality increases, the exponentially growing configuration space and the breakdown of full symmetry assumptions make existing approaches insufficient. These challenges highlight the need for novel paradigms to explore complex structures beyond human intuition.
@@
In this paper, we introduce PackingStar, a multi-agent reinforcement learning system that reformulates the Kissing Number Problem as a two-player matrix completion game in the Gram matrix domain. By operating directly on pairwise cosines rather than high-dimensional coordinates, this formulation eliminates numerical instability and computational inefficiency, enabling fully parallelizable GPU acceleration and large-scale sampling. Within PackingStar, the first agent incrementally fills matrix entries to construct larger and denser configurations. In early exploration stages, however, the candidate space is exponentially large and only local matrix information is observable, making erroneous placements inevitable. To address this challenge, the second agent performs delayed and selective corrections by leveraging a more global understanding of matrix, dynamically removing invalid or inconsistent entries and compressing the first agent’s search space. This interaction enables the joint system to converge toward larger configurations corresponding to higher kissing numbers. Meanwhile, the evolving matrices are iteratively decomposed into representative substructures that distill essential geometric patterns and impose structural constraints for subsequent games. Together, these mechanisms transform unconstrained combinatorial exploration into guided construction, substantially reducing reinforcement learning complexity in exponentially large spaces and enabling scalable discovery of high-dimensional configurations previously inaccessible to existing methods.
@@
Figure 1: AI breakthroughs in the Kissing Number Problem. Grey markers show the best previously known lower bounds. The blue triangle marks the 11D improvement achieved by AlphaEvolve (2025) [?]. The red stars indicate new lower bounds $K_{\mathrm{new}}(n)$ discovered by PackingStar, with the inset reporting the corresponding increments $\Delta K(n)=K_{\mathrm{new}}(n)-K_{\mathrm{prev}}(n)$ compared to the previous best $K_{\mathrm{prev}}(n)$.
@@
PackingStar demonstrates remarkable scalability, successfully reproducing all known kissing number configurations in dimensions up to 31. Figure 1 illustrates some of PackingStar’s breakthrough discoveries on this problem. From dimensions 25 to 31, PackingStar surpasses the lower bounds reported in 2016 and also exceeds the construction form introduced in 2011 [?,?], achieving what appears to be the optimal construction form along this path; the resulting construction form are summarized in Table 2. Notably, the breakthrough configuration discovered in 25 dimensions exhibits a clear geometric pattern that corresponds closely to the substructure of Leech lattice. Across multiple metrics, this configuration shows strong convergence, suggesting that it may represent the optimal 25-dimensional arrangement, although a rigorous mathematical proof is still lacking. In 13 dimensions, PackingStar surpasses the previously known rational structure with 1130 spheres from 1971 [?], reaching a kissing number of 1146. Rational structures are those in which all pairwise cosines between sphere centers are rational. The non-rational structure with 1154 spheres is obtained by slightly adjusting the coordinates of some spheres in the 1130-sphere configuration, later reached 1154 in 1999 [?]. Although the new 13-dimensional rational configuration does not exceed the non-rational record, it is of considerable mathematical value. Its fully rational inter-sphere cosines allow for exact analytic study without relying on numerical approximations. This offers insights into the effects of different constraints on kissing numbers and may inspire higher-density packings. Rational structures also facilitate more efficient encoding and decoding in information and communication applications, highlighting their practical significance alongside their theoretical and geometric value. Beyond these specific dimensions, PackingStar uncovers over 6000 new configurations in 14 dimensions and identifies additional configurations corresponding to current lower bounds in dimensions 12, 13, 15, and others. The sheer volume of these newly discovered arrangements provides rich opportunities for further investigation and may ultimately lead to configurations with even higher kissing numbers. In addition, PackingStar achieves significant new records for generalized kissing configurations under tighter angular constraints, including 12-dimensional, 20-dimensional and 21-dimensional configurations with 81, 405 and 567 spheres under the $1/4$ cosine constraint, as well as 14-dimensional and 17-dimensional configurations with 252 and 578 spheres under the $1/3$ constraint, respectively. Notably, several of the newly discovered configurations fundamentally challenge long-held antipodal paradigms. Rather than conforming to symmetry-driven constructions centered around antipodal pairing, these configurations exhibit asymmetric yet highly regular geometric patterns. In multiple cases, they reveal deep algebraic correspondences with finite simple groups and previously unnoticed geometric relationships across dimensions. These structural insights have, in turn, inspired mathematicians to extract new and more general construction principles. Building on patterns uncovered by PackingStar, subsequent human-designed configurations have further improved existing lower bounds in several dimensions, including 22 dimensions, illustrating a productive feedback loop between AI-driven discovery and mathematical theory. All of these results have been independently validated and are now included in the standard databases for kissing numbers and spherical codes. In summary, these results showcase AI’s power to understand high-dimensional geometry, reshape established mathematical intuitions, and drive progress on century-old geometry problems.
@@ Results
The second objective is to identify improved construction forms that assemble these subsets into full configurations. Both objectives are jointly addressed within the unified PackingStar framework, which simultaneously increases the achievable size of $S_{i}$ and uncovers superior construction forms. Our findings show that PackingStar not only discovers larger subsets but also identifies a better construction form; moreover, evidence from convergence and geometric consistency strongly suggests that the new template is optimal within this construction pathway. These advances yield especially pronounced improvements in dimensions 29 and 31. Table 2 reports additional new construction forms that deviate from the classical form [?,?]. In dimensions 26263030, the system consistently augments the Leech-lattice component by adding lower-dimensional parts $K(k)$ where k=n24k=n-24 (e.g., adding $K(2)$ in dimension $26$). In 31 dimensions, PackingStar uncovers a novel assembly pattern by partitioning a 7-dimensional kissing configuration into 42 (optimal) disjoint unit-radius equilateral triangles, resulting in an $84$-fold weighted $S_{i}$, surpassing the previous $75$-fold weighted $S_{i}$. Similarly, in 29 dimensions, PackingStar embeds 12 (optimal) disjoint unit-radius equilateral triangles into the 5-dimensional kissing configuration to generate a $26$-fold weighted $S_{i}$, exceeding the prior $24$-fold weighted $S_{i}$. These advances yield strictly larger kissing configurations in all dimensions from $25$ to $31$. The resulting new lower bounds $K_{\mathrm{new}}(n)$ are summarized in Table 2. Compared with the previous best values $K_{\mathrm{prev}}(n)$, our constructions increase the kissing number by $8,38,68,152,1224,456,$ and $5476$ spheres in dimensions $25$ through $31$, respectively. These improvements are also highlighted by the red markers in Fig. 1. Table 4 summarizes the cosine value among sphere pairs in the new configurations.
New records for generalized kissing numbers
Table 3: New maximal values $K_{\mathrm{new}}(n,\alpha)$ for generalized kissing configurations with maximal inner product at most $\alpha$, obtained by PackingStar, compared with the previous best known values $K_{\mathrm{prev}}(n,\alpha)$. Bold entries indicate improved configurations discovered by PackingStar, and $\star$ denotes human-improved constructions inspired by structures from PackingStar.
12 $1/4$ $\{-\tfrac{1}{2},\ -\tfrac{1}{8},\ \tfrac{1}{4}\}$ 81 79 [?]
20 $1/4$ $\{-\tfrac{1}{2},\ -\tfrac{1}{8},\ \tfrac{1}{4}\}$ 405 378 [?]
21 $1/4$ $\{-\tfrac{1}{2},\ -\tfrac{1}{8},\ \tfrac{1}{4}\}$ 567 554 [?]
22 $1/5$ $\{-1,\ -\tfrac{1}{5},\ \tfrac{1}{5}\}$ 352 338 [?]
As shown in Table 3, the new constructions from PackingStar yield new record for $K(n,\alpha)$ in several dimensions. In 12, 20 and 21 dimensions under the $1/4$ constraint, we obtain $K_{\mathrm{new}}(12,1/4)=81$, $K_{\mathrm{new}}(20,1/4)=405$ and $K_{\mathrm{new}}(21,1/4)=567$, surpassing the previous records of 79, 378 and 554, respectively. We also report improvements under the $1/3$ constraint: in dimensions 14 and 17, we obtain $K_{\mathrm{new}}(14,1/3)=252$ and $K_{\mathrm{new}}(17,1/3)=578$, which beat the prior best known records of 240 and 546.
Notably, the $K_{\mathrm{new}}(12,1/4)$ configuration contains no antipodal pairs; however, each of the 81 spheres has the same number of neighbors at each distance, forming a highly regular local structure. Such a combination of high local regularity without any antipodal pair appears to be rare in spherical code. We find several distinct variants of $K_{\mathrm{new}}(12,1/4)$ configuration, among which the most symmetric realization turns out to be the Kronecker product of the Schläfli spherical code with the unit equilateral triangle. This arrangement possesses a transitive symmetry group of order 311040. In dimension 17, PackingStar found a significant improvement: a configuration with 578 spheres, exceeding the previously known 546. All three geometrically non-isomorphic variants discovered share the same cosine set. Analysis of these variants reveals a remarkable structural property: the $K_{\mathrm{new}}(17,1/3)$ configuration can be interpreted as essentially two copies of the 16-dimensional 288-sphere configuration, which is known to be optimal under the cosine $1/4$ constraint, together with two additional points, i.e., $578=288\times 2+2$. This decomposition provides strong evidence that 578 is very likely optimal for the cosine $1/3$ constraint in dimension 17.